A 710 spacecraft has total energy -0.58 and is in circular orbit around the Sun. Find (a) its orbital radius, (b) its kinetic energy, and (c) its speed
1
Expert's answer
2012-06-19T09:13:28-0400
E=Ek+Ep
Ek=(mv^2)/2
Ep=mgR
m=710
E=0.58
a)
E=mgR+(mv^2)/2
R=((mv^2)/2-E)/mg
b)
Ek=E-Ep=E-mgR
c)
E=mgR+(mv^2)/2
v=sqrt(2(E-mgR)/m^2)